# Recursive variational Gaussian approximation with the Whittle likelihood for linear non-Gaussian state space models

### v2

@@ Abstract
Parameter inference for linear and non-Gaussian state space models is challenging because the likelihood function contains an intractable integral over the latent state variables. While Markov chain Monte Carlo (MCMC) methods provide exact samples from the posterior distribution as the number of samples go to infinity, they tend to have high computational cost, particularly for observations of a long time series. Variational Bayes (VB) methods are a useful alternative when inference with MCMC methods is computationally expensive. VB methods approximate the posterior density of the parameters by a simple and tractable distribution found through optimisation. In this paper, we propose a novel sequential variational Bayes approach that makes use of the Whittle likelihood for computationally efficient parameter inference in this class of state space models. Our algorithm, which we call Recursive Variational Gaussian Approximation with the Whittle Likelihood (R-VGA-Whittle), updates the variational parameters by processing data in the frequency domain. At each iteration, R-VGA-Whittle requires the gradient and Hessian of the Whittle log-likelihood, which are available in closed form for a wide class of models. Through several examples using a linear Gaussian state space model and a univariate/bivariate non-Gaussian stochastic volatility model, we show that R-VGA-Whittle provides good approximations to posterior distributions of the parameters and is very computationally efficient when compared to asymptotically exact methods such as Hamiltonian Monte Carlo.
Keywords: Autoregressive process; Fast Fourier transform; Hamiltonian Monte Carlo; Stationary time series; Stochastic volatility models.
@@ 1 Introduction
There are two main classes of VB methods: batch and sequential. A batch VB framework (e.g., Tran et al.,, 2017; Tan and Nott,, 2018; Loaiza-Maya et al.,, 2022; Gunawan et al.,, 2024) requires an optimisation algorithm to repeatedly process the entire dataset until variational parameters converge to a fixed point. A sequential VB framework, on the other hand, updates variational parameters using one observation at a time. Sequential methods only pass through the entire dataset once, and are thus ideal for settings with large datasets. Sequential VB methods are available for statistical models for independent data (Lambert et al.,, 2022), for classical autoregressive models for time-series data (Tomasetti et al.,, 2022), for generalised linear mixed models (Vu et al.,, 2024), and for SSMs where the state evolution model is linear and Gaussian (Beal,, 2003, Chapter 5; Zammit Mangion et al.,, 2011; Mangion et al.,, 2011). Campbell et al., (2021) proposed a sequential VB approach that is applicable to an SSM with non-linear state dynamics and non-Gaussian measurement errors, but this approach can be computationally expensive as it requires solving an optimisation problem and a regression task for each time step.
@@
We demonstrate the performance of R-VGA-Whittle in simulation studies involving a linear Gaussian state space model and univariate and bivariate non-Gaussian stochastic volatility models. We show that R-VGA-Whittle is accurate and is much faster than HMC with the exact likelihood as well as HMC based on the Whittle likelihood; hereafter, we refer to these two methods as HMC-exact and HMC-Whittle, respectively. We subsequently apply R-VGA-Whittle to analyse exchange rate data.
@@ 2.1 State space models
$\mathbf{X}_{t}\mid(\mathbf{X}_{t-1}=\mathbf{x}_{t-1})\sim p(\mathbf{x}_{t}\mid\mathbf{x}_{t-1},\bm{\theta}),\quad t=2,\dots,T,$ (1)
where $\bm{\theta}$ is a vector of model parameters. The observations $\{\mathbf{y}_{t}\}_{t=1}^{T}$ are related to the latent states via
$\mathbf{Y}_{t}\mid(\mathbf{X}_{t}=\mathbf{x}_{t})\sim p(\mathbf{y}_{t}\mid\mathbf{x}_{t},\bm{\theta}),\quad t=1,\dots,T,$ (2)
where each $\mathbf{Y}_{t}$ is assumed to be conditionally independent of the other observations, when conditioned on the latent state $\mathbf{x}_{t}$. A linear SSM is one where $\mathbf{X}_{t}$ and $\mathbf{Y}_{t}$ in (1) and (2) can each be expressed as the sum of two components: a linear function of their conditioning variables and an exogenous random component.
@@
The Whittle likelihood (Whittle,, 1953) is a frequency-domain approximation of the exact likelihood, and offers a computationally efficient alternative for parameter inference in SSMs. Transforming the data to the frequency domain “bypasses” the need for integrating out the states when computing the likelihood, as the Whittle likelihood only depends on the model parameters and not the latent variables. In the following sections, we review the Whittle likelihood and its properties, and then propose a sequential VB approach that uses the Whittle likelihood for parameter inference.
@@ 2.2 The Whittle likelihood
The Whittle likelihood has been widely used to estimate the parameters of complex time series or spatial models. It is more computationally efficient to evaluate than the exact likelihood, although it is approximate (e.g., Sykulski et al.,, 2019). Some applications of the Whittle likelihood include parameter estimation with long memory processes (e.g., Robinson,, 1995; Hurvich and Chen,, 2000; Giraitis and Robinson,, 2001; Abadir et al.,, 2007; Shao and Wu, 2007b, ; Giraitis et al.,, 2012), and with spatial models with irregularly spaced data (e.g., Fuentes,, 2007; Matsuda and Yajima,, 2009). The Whittle likelihood is also used for estimating parameters of vector autoregressive moving average (VARMA) models with heavy-tailed errors (She et al.,, 2022), and of vector autoregressive tempered fractionally integrated moving average (VARTFIMA) models (Villani et al.,, 2022).
@@ 2.2.1 Univariate Whittle likelihood
Here, the Whittle likelihood is evaluated by summing over the non-negative frequencies only, since $f(\omega_{k};\bm{\theta})$ and $I(\omega_{k})$ are both symmetric about the origin for real-valued time series. The term $\omega_{k}=0$ is not included since $I(0)=0$ when the time series is de-meaned. The term $\omega_{k}=\pi$ is also omitted since it has a different distribution from the other frequencies and its influence is asymptotically negligible (Salomone et al.,, 2020; Villani et al.,, 2022).
Asymptotically, it can be shown that the parameter estimates obtained by maximising the exact likelihood and those obtained by maximising the Whittle likelihood are equivalent (Dzhaparidze and Yaglom,, 1983). The calculation of the Whittle likelihood involves evaluation of the periodogram, which does not depend on the parameters $\bm{\theta}$ and can, therefore, be pre-computed at a cost of $O(T\log T)$ via the Fast Fourier Transform (FFT, Cooley and Tukey,, 1965). Since the periodogram ordinates are asymptotically independent (Whittle,, 1953), the logarithm of the Whittle likelihood is a sum of the individual log-likelihood terms computed over different frequencies.
@@ 2.2.2 Multivariate Whittle likelihood
The Whittle likelihood for the multivariate series $\{\mathbf{y}_{t}\}_{t=1}^{T}$ is (Villani et al.,, 2022):
$l_{W}(\bm{\theta})=-\sum_{k=1}^{\lfloor\frac{T-1}{2}\rfloor}\left(\log\lvert\mathbf{f}(\omega_{k};\bm{\theta})\rvert+\operatorname{Tr}(\mathbf{f}(\omega_{k};\bm{\theta}))^{-1}\mathbf{I}(\omega_{k})\right).$ (10)
@@ 2.3 R-VGA-Whittle
At each iteration ${i=1,\dots,N}$, the R-VGA algorithm requires the gradient and Hessian of the log-likelihood contribution from each observation, $\log p(\mathbf{y}_{i}\mid\bm{\theta})$. In a setting with independent observations, these likelihood contributions are easy to obtain; however, in SSMs where the latent states follow a Markov process and are hence temporally dependent, the log-likelihood function is intractable, as shown in Section 2.1. The gradient and Hessian of $\log p(\mathbf{y}_{t}\mid\bm{\theta})$, ${t=1,\dots,T}$, are hence also intractable. While these quantities can be estimated via particle filtering (e.g., Doucet and Shephard,, 2012), one would need to run a particle filter from the first time period to the current time period for every step in the sequential algorithm. This quickly becomes infeasible for long time series.
@@
We now extend the R-VGA algorithm of Lambert et al., (2022) to state space models by replacing the time domain likelihood with a frequency domain alternative: the Whittle likelihood. From (6) and (10) we see that the log of the Whittle likelihood is a sum over individual log-likelihood components, a consequence of the asymptotic independence of the periodogram frequency components (Shao and Wu, 2007a, ). We replace the quantities ${\nabla_{\bm{\theta}}\log p(\mathbf{y}_{i}\mid\bm{\theta})}$ and ${\nabla_{\bm{\theta}}^{2}\log p(\mathbf{y}_{i}\mid\bm{\theta})}$ in Algorithm 1 with the gradient and Hessian of the Whittle log-likelihood evaluated at each frequency. In the univariate case, the individual log-likelihood contribution at frequency $\omega_{k}$ is
$l_{W_{k}}(\bm{\theta})=\log f(\omega_{k};\bm{\theta})+\frac{I(\omega_{k})}{f(\omega_{k};\bm{\theta})},\quad k=1,\dots,\left\lfloor\frac{T-1}{2}\right\rfloor,$ (11)
@@
$l_{W_{k}}(\bm{\theta})=\log\lvert\mathbf{f}(\omega_{k};\bm{\theta})\rvert+\operatorname{Tr}(\mathbf{f}(\omega_{k};\bm{\theta}))^{-1}\mathbf{I}(\omega_{k}),\quad k=1,\dots,\left\lfloor\frac{T-1}{2}\right\rfloor,$ (12)
for $\omega_{k}=\frac{2\pi k}{T}$. The log-likelihood contributions $l_{W_{k}}(\bm{\theta})$ are available for models where the power spectral density is available analytically, such as models where the latent states follow a VARMA (She et al.,, 2022) or a VARTFIMA (Villani et al.,, 2022) process. The gradient ${\nabla_{\bm{\theta}}l_{W_{k}}(\bm{\theta})}$ and Hessian ${\nabla_{\bm{\theta}}^{2}l_{W_{k}}(\bm{\theta})}$ are therefore also available in closed form for these models. At each VB update, ${\nabla_{\bm{\theta}}l_{W_{k}}(\bm{\theta})}$ and ${\nabla_{\bm{\theta}}^{2}l_{W_{k}}(\bm{\theta})}$ are used to update the variational parameters for $k=1,\dots,\left\lfloor\frac{T-1}{2}\right\rfloor$. Their expectations with respect to $q_{k-1}(\bm{\theta})$ are approximated using Monte Carlo samples ${\bm{\theta}^{(s)}\sim q_{k-1}(\bm{\theta})}$, ${s=1,\dots,S}$:
@@
         Set $a=1/D$, $\bm{\mu}_{k,0}=\bm{\mu}_{k-1},\mathbf{\Sigma}_{k,0}=\mathbf{\Sigma}_{k-1}$
@@
              Set $q_{k,l-1}(\bm{\theta})=\mathrm{Gau}(\bm{\mu}_{k,l-1},\mathbf{\Sigma}_{k,l-1})$.
              $\bm{\mu}_{k,l}=\bm{\mu}_{k,l-1}+a\mathbf{\Sigma}_{k,l}\mathbb{E}_{q_{k,l-1}}(\nabla_{\bm{\theta}}l_{W_{k}}(\bm{\theta}))$
              $\mathbf{\Sigma}_{k,l}^{-1}=\mathbf{\Sigma}_{k,l-1}^{-1}-a\mathbb{E}_{q_{k,l-1}}(\nabla_{\bm{\theta}}^{2}l_{W_{k}}(\bm{\theta}))$
@@
         $\bm{\mu}_{k}=\bm{\mu}_{k-1}+\mathbf{\Sigma}_{k}\mathbb{E}_{q_{k-1}}({\nabla_{\bm{\theta}}l_{W_{k}}(\bm{\theta})})$
         $\mathbf{\Sigma}_{k}^{-1}=\mathbf{\Sigma}_{k-1}^{-1}-\mathbb{E}_{q_{k-1}}({\nabla_{\bm{\theta}}^{2}l_{W_{k}}(\bm{\theta})})$
@@ 2.4 R-VGA-Whittle with block updates
The spectral density of a process captures the distribution of signal power (variance) across frequency (Percival and Walden,, 1993). The periodogram is an estimate of the spectral density. In Figure 1(a), we show the periodogram for data generated from the linear Gaussian SSM that we consider in Section 3.1. The variance (power) of the signal gradually decreases as the frequency increases, which means that lower frequencies contribute more to the total power of the process. A similar pattern is seen in Figure 1(b), which shows the periodogram for data generated from a univariate stochastic volatility model that we consider in Section 3.2. In this case there is an even larger concentration of power at the lower frequencies.
@@
Figure 1: Plot of raw periodograms of the data (black) and smoothed periodograms (light red) for two univariate models specified in Section 3. Vertical red dashed lines show the frequency cutoffs at half-power (3dB). Frequencies after this cutoff are processed in blocks.
@@
Figure 2: Plot of raw periodograms (black) and smoothed periodograms (light red) for the bivariate stochastic volatility model specified in Section 3. Vertical dashed lines show the 3dB frequency cutoff points for each series. In this bivariate case, the cutoff point we use for R-VGA-Whittle is the higher of the two.
Frequencies that carry less power have less effect on R-VGA-Whittle updates. To speed up the algorithm and reduce the number of gradient/Hessian calculations, we therefore process angular frequency contributions with low power, namely those beyond the 3dB cutoff point, together in “blocks”. We define the 3dB cutoff point as the angular frequency at which the power drops to half the maximum power in the (smoothed) periodogram, which we obtain using Welch’s method (Welch,, 1967). Section S3 studies several cutoff points and empirically justifies our choice of a 3dB cutoff.
@@
that is, by the sum of the Whittle log-likelihood contributions of the frequencies in the block.
For a multivariate time series, we compute the periodogram for each individual time series along with its corresponding 3dB cutoff point, and then use the highest cutoff point in R-VGA-Whittle. An example of bivariate time series periodograms along with their 3dB cutoffs points is shown in Figure 2.
We summarise R-VGA-Whittle with block updates in Algorithm 3.
@@
Input: observations $\mathbf{y}_{1},\dots,\mathbf{y}_{T}$, initial values $\bm{\mu}_{0}$ and $\mathbf{\Sigma}_{0}$, number of angular frequencies to damp $n_{damp}$, number of damping steps $D$, number of angular frequencies to update individually $\tilde{n}>n_{damp}$, block length $B$.
@@
     Update $\bm{\mu}_{\tilde{k}},\mathbf{\Sigma}_{\tilde{k}}$ as in Algorithm 2
@@
     Set $\tilde{k}=\tilde{n}+b$
     $\bm{\mu}_{\tilde{k}}=\bm{\mu}_{{\tilde{k}}-1}+\mathbf{\Sigma}_{\tilde{k}}\mathbb{E}_{q_{{\tilde{k}}-1}}(\nabla_{\bm{\theta}}l_{W}^{b}(\bm{\theta}))$, where $l_{W}^{b}(\bm{\theta})$ is given in (13)
     $\mathbf{\Sigma}_{\tilde{k}}^{-1}=\mathbf{\Sigma}_{{\tilde{k}}-1}^{-1}-\mathbb{E}_{q_{{\tilde{k}}-1}}(\nabla_{\bm{\theta}}^{2}l_{W}^{b}(\bm{\theta}))$

### v3

# output-myclipboard

@@ Abstract
Parameter inference for linear and non-Gaussian state space models is challenging because the likelihood function contains an intractable integral over the latent state variables. While Markov chain Monte Carlo (MCMC) methods provide exact samples from the posterior distribution as the number of samples goes to infinity, they tend to have high computational cost, particularly for observations of a long time series. When inference with MCMC methods is computationally expensive, variational Bayes (VB) methods are a useful alternative. VB methods approximate the posterior density of the parameters with a simple and tractable distribution found through optimisation. This work proposes a novel sequential VB approach that makes use of the Whittle likelihood for computationally efficient parameter inference in linear, non-Gaussian state space models. Our algorithm, called Recursive Variational Gaussian Approximation with the Whittle Likelihood (R-VGA-Whittle), updates the variational parameters by processing data in the frequency domain. At each iteration, R-VGA-Whittle requires the gradient and Hessian of the Whittle log-likelihood, which are available in closed form. Through several examples involving a linear Gaussian state space model; a univariate/bivariate stochastic volatility model; and a state space model with Student’s t measurement error, where the latent states follow an autoregressive fractionally integrated moving average (ARFIMA) model, we show that R-VGA-Whittle provides good approximations to posterior distributions of the parameters, and that it is very computationally efficient when compared to asymptotically exact methods such as Hamiltonian Monte Carlo.
Keywords: ARFIMA process; Fast Fourier transform; Hamiltonian Monte Carlo; Stationary time series; Stochastic volatility models.
@@ 1 Introduction
There are two main classes of VB methods: batch and sequential. A batch VB framework (e.g., Tran et al.,, 2017; Tan and Nott,, 2018; Loaiza-Maya et al.,, 2022; Gunawan et al.,, 2024) requires an optimisation algorithm to repeatedly process the entire dataset until variational parameters converge to a fixed point. A sequential VB framework, on the other hand, updates variational parameters using one observation at a time. Sequential methods only pass through the entire dataset once, and are thus ideal for use with large datasets. Sequential VB methods are available for statistical models for independent data (Lambert et al.,, 2022), for classical autoregressive models for time-series data (Tomasetti et al.,, 2022), for generalised linear mixed models (Vu et al.,, 2024), and for SSMs where the state evolution model is linear and Gaussian (Beal,, 2003, Chapter 5; Zammit Mangion et al.,, 2011). Campbell et al., (2021) proposed a sequential VB approach that is applicable to an SSM with non-linear state dynamics and non-Gaussian measurement errors, but this approach can be computationally expensive as it requires solving an optimisation problem and a regression task at each time step.
@@
We demonstrate the performance of R-VGA-Whittle in simulation studies involving a linear Gaussian state space model, univariate and bivariate stochastic volatility models, and a state space model with measurement errors that follow a Student’s $t$ distribution and latent states that follow an autoregressive fractionally integrated moving average model (ARFIMA; Granger and Joyeux, (1980)). We show that R-VGA-Whittle is accurate and is often much faster than both HMC with the exact likelihood and HMC based on the Whittle likelihood; hereafter, we refer to these two methods as HMC-exact and HMC-Whittle, respectively. We subsequently apply R-VGA-Whittle to analyse exchange rate data.
@@ 2.1 State space models
$\mathbf{x}_{t}\mid\mathbf{x}_{t-1},\bm{\theta}\sim p(\mathbf{x}_{t}\mid\mathbf{x}_{t-1},\bm{\theta}),\quad t=2,\dots,T,$ (1)
where $\bm{\theta}$ is a vector of model parameters. The observations $\{\mathbf{y}_{t}\}_{t=1}^{T}$ are related to the latent states via the data model
$\mathbf{y}_{t}\mid\mathbf{x}_{t},\bm{\theta}\sim p(\mathbf{y}_{t}\mid\mathbf{x}_{t},\bm{\theta}),\quad t=1,\dots,T,$ (2)
[…]
$\mathbf{y}_{t}=\mathbf{x}_{t}+\bm{\varepsilon}_{t},\quad t=1,\dots,T,$ (3)
where $\mathbf{x}_{t}$ is the $d$-dimensional latent state variables and $\bm{\varepsilon}_{t}$ is a (possibly non-Gaussian) vector-valued noise process, both of which have closed-form spectral densities defined in Section 2.2.
@@
[…]
@@ 2.2 The Whittle likelihood
The Whittle likelihood has been widely used to estimate the parameters of complex time series or spatial models. It is more computationally efficient to evaluate than the exact likelihood, although it is approximate (e.g., Sykulski et al.,, 2019). Some applications of the Whittle likelihood include parameter estimation with long memory processes (e.g., Robinson,, 1995; Hurvich and Chen,, 2000; Giraitis and Robinson,, 2001; Abadir et al.,, 2007; Shao and Wu, 2007b, ; Giraitis et al.,, 2012), and with spatial models with irregularly spaced data (e.g., Fuentes,, 2007; Matsuda and Yajima,, 2009). The Whittle likelihood is also used for estimating parameters of vector autoregressive moving average (VARMA) models with heavy-tailed errors (She et al.,, 2022), and of vector autoregressive tempered fractionally integrated moving average (VARTFIMA) models (Villani et al.,, 2024).
@@ 2.2.1 Univariate Whittle likelihood
Here, the Whittle likelihood is evaluated by summing over the non-negative frequencies only, since $f(\omega_{k};\bm{\theta})$ and $I(\omega_{k})$ are both symmetric about the origin for real-valued time series. The term $\omega_{k}=0$ is not included since $I(0)=0$ when the time series is de-meaned. The term $\omega_{k}=\pi$ is also omitted since it has a different distribution from the other frequencies and its influence is asymptotically negligible (Salomone et al.,, 2020; Villani et al.,, 2024).
As $T$ tends to infinity, it can be shown that the parameter estimates obtained by maximising the exact Gaussian likelihood and those obtained by maximising the Whittle likelihood are equivalent (Dzhaparidze and Yaglom,, 1983). The calculation of the Whittle likelihood involves evaluation of the periodogram, which does not depend on the parameters $\bm{\theta}$ and can, therefore, be pre-computed at a cost of $O(T\log T)$ via the Fast Fourier Transform (FFT, Cooley and Tukey,, 1965). Since the periodogram ordinates are asymptotically independent (Shao and Wu, 2007a, ), the logarithm of the Whittle likelihood is a sum of the individual log-likelihood terms computed over different frequencies.
@@ 2.2.2 Multivariate Whittle likelihood
The Whittle likelihood for the multivariate series $\{\mathbf{y}_{t}\}_{t=1}^{T}$ is (Villani et al.,, 2024):
$l_{W}(\bm{\theta})=-\sum_{k=1}^{\lfloor\frac{T-1}{2}\rfloor}\left(\log\lvert\mathbf{f}(\omega_{k};\bm{\theta})\rvert+\operatorname{Tr}(\mathbf{f}(\omega_{k};\bm{\theta})^{-1}\mathbf{I}(\omega_{k}))\right).$ (11)
@@ 2.3 R-VGA-Whittle
At each iteration ${i=1,\dots,N}$, the R-VGA algorithm requires the gradient and Hessian of the log-likelihood contribution from each observation, $\log p(\mathbf{y}_{i}\mid\bm{\theta})$. In a setting with independent observations, these log-likelihood contributions are easy to obtain; however, in SSMs where the latent states follow a Markov process and are hence temporally dependent, the log-likelihood function is intractable, as shown in Section 2.1. The gradient and Hessian of $\log p(\mathbf{y}_{t}\mid\bm{\theta})$, ${t=1,\dots,T}$, are hence also intractable. While these quantities can be estimated via particle filtering (e.g., Doucet and Shephard,, 2012), one would need to run a particle filter from the first time period to the current time period for every step in the sequential algorithm. This quickly becomes infeasible for long time series.
@@
We now extend the R-VGA algorithm of Lambert et al., (2022) to state space models by replacing the time domain likelihood with a frequency domain alternative: the Whittle likelihood. From (7) and (11) we see that the log of the Whittle likelihood is a sum over individual log-likelihood components, a consequence of the asymptotic independence of the periodogram frequency components. We replace the quantities ${\nabla_{\bm{\theta}}\log p(\mathbf{y}_{i}\mid\bm{\theta})}$ and ${\nabla_{\bm{\theta}}^{2}\log p(\mathbf{y}_{i}\mid\bm{\theta})}$ in Algorithm 1 with the gradient and Hessian of the Whittle log-likelihood evaluated at each frequency. In the univariate case, the individual log-likelihood contribution at frequency $\omega_{k}$ is
$l_{W_{k}}(\bm{\theta})=-\log f(\omega_{k};\bm{\theta})-\frac{I(\omega_{k})}{f(\omega_{k};\bm{\theta})},\quad k=1,\dots,\left\lfloor\frac{T-1}{2}\right\rfloor,$ (12)
@@
$l_{W_{k}}(\bm{\theta})=-\log\lvert\mathbf{f}(\omega_{k};\bm{\theta})\rvert-\operatorname{Tr}(\mathbf{f}(\omega_{k};\bm{\theta})^{-1}\mathbf{I}(\omega_{k})),\quad k=1,\dots,\left\lfloor\frac{T-1}{2}\right\rfloor,$ (13)
for $\omega_{k}=\frac{2\pi k}{T}$. The log-likelihood contributions $l_{W_{k}}(\bm{\theta})$ are available for models where the power spectral density is available analytically, such as models where the latent states follow a VARMA (She et al.,, 2022) or a VARTFIMA (Villani et al.,, 2024) process. The gradient ${\nabla_{\bm{\theta}}l_{W_{k}}(\bm{\theta})}$ and Hessian ${\nabla_{\bm{\theta}}^{2}l_{W_{k}}(\bm{\theta})}$ are therefore also available in closed form for these models. At each VB update, ${\nabla_{\bm{\theta}}l_{W_{k}}(\bm{\theta})}$ and ${\nabla_{\bm{\theta}}^{2}l_{W_{k}}(\bm{\theta})}$ are used to update the variational parameters for $k=1,\dots,\left\lfloor\frac{T-1}{2}\right\rfloor$. Their expectations with respect to $q_{k-1}(\bm{\theta})$ are approximated using Monte Carlo samples ${\bm{\theta}^{(s)}\sim q_{k-1}(\bm{\theta})}$, ${s=1,\dots,S}$:
@@
         Set $a=1/D$, $\bm{\mu}_{k,0}=\bm{\mu}_{k-1},\mathbf{\Sigma}_{k,0}=\mathbf{\Sigma}_{k-1}$.
@@
              Set $q_{k,l-1}(\bm{\theta})=\mathrm{Gau}(\bm{\mu}_{k,l-1},\mathbf{\Sigma}_{k,l-1})$,
              $\bm{\mu}_{k,l}=\bm{\mu}_{k,l-1}+a\mathbf{\Sigma}_{k,l}\mathbb{E}_{q_{k,l-1}}(\nabla_{\bm{\theta}}l_{W_{k}}(\bm{\theta}))$,
              $\mathbf{\Sigma}_{k,l}^{-1}=\mathbf{\Sigma}_{k,l-1}^{-1}-a\mathbb{E}_{q_{k,l-1}}(\nabla_{\bm{\theta}}^{2}l_{W_{k}}(\bm{\theta}))$.
@@
         $\bm{\mu}_{k}=\bm{\mu}_{k-1}+\mathbf{\Sigma}_{k}\mathbb{E}_{q_{k-1}}({\nabla_{\bm{\theta}}l_{W_{k}}(\bm{\theta})})$,
         $\mathbf{\Sigma}_{k}^{-1}=\mathbf{\Sigma}_{k-1}^{-1}-\mathbb{E}_{q_{k-1}}({\nabla_{\bm{\theta}}^{2}l_{W_{k}}(\bm{\theta})})$.
@@ 2.4 R-VGA-Whittle with block updates
[…]
@@
Figure 1: Plot of raw periodograms of the data (black) and smoothed periodograms (light red) for the univariate linear Gaussian state space model of Section 3.1 and the univariate stochastic volatility of Section 3.2. Vertical red dashed lines show the frequency cutoffs at half-power (3dB). Frequencies after this cutoff are processed in blocks.
@@
Figure 2: Plot of raw periodograms (black) and smoothed periodograms (light red) for the bivariate stochastic volatility model of Section 3.3. Vertical dashed lines show the 3dB frequency cutoff points for each series. In this bivariate case, the cutoff point we use for R-VGA-Whittle is the higher of the two.
@@
that is, by the sum of the Whittle log-likelihood contributions of the frequencies in the block. For a multivariate time series, we compute the periodogram for each individual time series along with its corresponding 3dB cutoff point, and then use the highest cutoff point in R-VGA-Whittle. An example of bivariate time series periodograms along with their 3dB cutoffs points is shown in Figure 2.
We summarise the R-VGA-Whittle algorithm with block updates in Algorithm 3, under the assumption that there are no high-frequency components with substantial power beyond $\tilde{n}$.
@@
Input: observations $\mathbf{y}_{1},\dots,\mathbf{y}_{T}$, initial values $\bm{\mu}_{0}$ and $\mathbf{\Sigma}_{0}$, number of angular frequencies to damp $n_{damp}$, number of damping steps $D$, number of angular frequencies to update individually $\tilde{n}>n_{damp}$, block length $B$. Assume that there are no high-frequency components with substantial power beyond $\tilde{n}$.
@@
     Update $\bm{\mu}_{\tilde{k}},\mathbf{\Sigma}_{\tilde{k}}$ as in Algorithm 2.
@@
     Set $\tilde{k}=\tilde{n}+b$,
     $\bm{\mu}_{\tilde{k}}=\bm{\mu}_{{\tilde{k}}-1}+\mathbf{\Sigma}_{\tilde{k}}\mathbb{E}_{q_{{\tilde{k}}-1}}(\nabla_{\bm{\theta}}l_{W}^{b}(\bm{\theta}))$, where $l_{W}^{b}(\bm{\theta})$ is given in (14),
     $\mathbf{\Sigma}_{\tilde{k}}^{-1}=\mathbf{\Sigma}_{{\tilde{k}}-1}^{-1}-\mathbb{E}_{q_{{\tilde{k}}-1}}(\nabla_{\bm{\theta}}^{2}l_{W}^{b}(\bm{\theta}))$.